Maxx Industry is regularly supplying some critical parts to your assembly line. These parts are either good or defective. During each delivery a sample of 500 units is sampled and the number of defective units in the sample is noted. The average defect rate is 3.5%.

Develop the upper and lower control limits (z=3).


1
Expert's answer
2021-11-24T02:19:02-0500

Considering the given situation for calculating the UCL and LCL, We can identify the control limit using the C-chart. In this case the defective rate is given in counting percentage which means, we can apply the C-chart method to find the control limits. Based on the information, the calculation will be like given below

"UCL=\bar{c}+3\sqrt{\bar{c}}"

"LCL=\bar{c}-3\sqrt{\bar{c}}"

"Here, \bar{c}=3.5\%of 500=17.5"

"UCL=17.5+3\sqrt{17.5}=17.5+12.5=30units"

"LCL=17.5-3\sqrt{17.5}=17.5-12.5=5units"

Based on the calculation, we identified the UCL as 30 unit and LCL as 5 unit using the C-chart method of control chart.